OpenStudy (anonymous):

What is the distance between P(–4, 3) and Q(6, 1)? Round to the nearest tenth

OpenStudy (anonymous):

Use the slope formula y2-y1/x2-x1 1-3/6+4 = -2/10 = -1/5 = -0.2 Does that make sense?

OpenStudy (anonymous):

no im sorry. im really bad at this stuff

OpenStudy (campbell_st):

well thats the slope and you need to find distance

OpenStudy (campbell_st):

and you forgot to square the brackets

OpenStudy (anonymous):

Good call again Oh good call campbell Distance formula is sqr[(x2-x1)^2 + (y2-y1)^2] You can plug in the numbers from there.